Find a counter example to the statement that every even positive integer can be written as the sum of the squares of three integers.


1
Expert's answer
2020-11-08T18:51:19-0500

Consider the integer number "n=28". Let us show that "n" can not be written as the sum of the squares of three integers. For this prove that the equation "a+b+c=28" has no solution, where "a,b,c" are squares of integers.


It followst that "a\geq0, b\geq 0, c\geq 0", and therefore "a\leq 28, b\leq 28, c\leq 28". Taking into account that "a,b,c" are squares of integers, we conclude that "a,b,c\in X=\{0,1,4,9,16,25\}".


Suppose that "a=25". Then "b+c=3", and we see that there are no "b,c\in X" such that "b+c=3". By analogy for the cases "b=25" and "c=25".


Suppose that "a=16". Then "b+c=12", and we see that there are no "b,c\in X" such that "b+c=12". By analogy for "b=16" and "c=16".


If "a\leq 9, b\leq 9, c\leq 9", then "a+b+c\leq 27", and there are no "a,b,c\in X" such that "a+b+c=28".


Therefore, "n=28" can not be written as the sum of the squares of three integers.